# FM-G-CAM: A Holistic Approach for Explainable AI in Computer Vision

### v1

@@ 3.3 Choosing an optimal value for (K)
Unlike other saliency map generation methods, FM-G-CAM utilises multiple classes to generate the final saliency map. We recommend choosing the top predicted classes to be used. The top 3 or top 5 classes are recommended to be used with FM-G-CAM even though the algorithm allows any number of arbitrarily chosen classes to be used in the saliency map generation. However, this can be different and will depend mainly based below three factors: 1) The nature of the use case, 2) Average area of interest on the image for each predictive class, 3) Average number of distinct predictive labels per image.
@@ 3.3.1 Importance of L2 Normalisation
As shown in Figure 3, L2-Normalisation across the class gradient maps makes FM-G-CAM sensitive and unbiased, especially in the presence of classes with relatively lower weighted activations, while remaining accurate based on the content of the image. Also, as explained in the algorithm 1, normalisation is applied only after the filtering process. This allows filtering to be done based on raw gradient-weighted activations, and independent of the normalisation process. L2 normalisation also aids in the reduction of unwanted noise, as shown in figure 2. This is due to the fact that it is applied across the saliency maps by comparing the values against each other.
@@ 3.3.2 Choice of Colour Maps
Based on the findings presented in Section 2.3, we recommend using a single-colour colour map for each class when K in Section 3.3 is set to the recommended value of four. This will allow proper visualisation of the CNN’s decision-making process without overly complicating the saliency map. We comparatively test these in section 5 with existing methods.
@@ 4 Implementation: XAI Inference Engine
For convenient use of FM-G-CAM, a Python library is available to the community [^1] https://pypi.org/project/xai-inference-engine. The implementation is based on the algorithm 1. The code is open source and available on GitHub [^2] https://github.com/SuienS/xai-inference-engine. The library allows the user to use their trained PyTorch CNN-based models with our library. The library will output the prediction results along with the generated FM-G-CAM saliency map explaining the results. Figure 4 shows the overview of the XAI Inference Engine.
@@
The library also allows changing the end activation function used in the saliency map generation to ReLU, ELU, or GeLU. Figure 5 shows the comparative results of the use of the different activation functions available. It shows that GeLU and ELU tend to attract more noise than ReLU. However, at the same time, those can be useful when we want the saliency maps to be more sensitive. In contrast, Grad-CAM remains identical across all activation functions.
@@ 5 Use cases: A Comparative Practical Guide for FM-G-CAM
In this section, we compare our method with the existing methods in generative saliency maps via a small subset of use cases. It is worth noting that the method we propose is characteristically different from virtually all of the existing methods due to the nature of considering multiple classes for the saliency map. To be as fair as possible, we perform Grad-CAM on all the classes that FM-G-CAM uses to generate the saliency map, as can be seen in the following sections. Also, we purposefully compare our technique only with Grad-CAM since any other improved or advanced version of Grad-CAM can be used within FM-G-CAM in the generation of the saliency map. We did not intend to improve the underlying theory of gradient-weighted activations for saliency map generation but to utilise that idea in a way to provide a more holistic explanation of CNN predictions. The following sections describe some example use cases of FM-G-CAM in comparison to Grad-CAM.
@@ 5.2 Medical AI
Another very interesting and important area where the proposed approach can be utilised is in medical AI. According to the literature, CNNs are a popular approach in the field of medical AI for a variety of diagnosis modalities [42, 43, 44, 45]. Therefore, an application of FM-G-CAM is explored in the diagnosis of chest X-Ray images. To this end, TorchXRayVision[46] weights are used for the DenseNet model[47]. Figure 7 shows explained outputs when classifying images selected from the validation set. The samples show that the FM-G-CAM presents a holistic and singular saliency map that can replace all other class-based Grad-CAM saliency maps. Further, FM-G-CAM comparatively visualises the saliency maps of the selected classes, portraying their relative importance, whereas Grad-CAM considers each class in isolation, completely neglecting the fact that they are all part of the same prediction result.
@@ 6 Discussion
Using the CNN activations to explain their prediction was first introduced by [30], based on previous work investigating the capabilities of CNNs in object detection [48]. These works formed a basis for the idea of Grad-CAM [10]. However, these studies only consider a target class and, by doing so, disregard a considerable portion of information regarding the CNN classification process. This study proposes that the activations should be calculated for the top-k classes, weighing them accordingly using their gradients, normalising them for fair representation, and finally fusing them. Hence, while Grad-CAM and FM-G-CAM are based on the aforementioned studies, they are characteristically and mathematically different.
The previously discussed Figure 6 shows that the single-class Grad-CAM approach is not always useful in situations where a bigger-picture approach is needed. Instead, it only shows why a specified class has its related prediction score. In most cases, this class is selected as the top predicted class. However, there is a significant difference in the majority of CNN-based predictor models in top-1444https://paperswithcode.com/sota/image-classification-on-imagenet accuracy vs top-$5$https://paperswithcode.com/sota/image-classification-on-imagenet?metric=Top%205%20Accuracy accuracy. The difference between the accuracy of top-1 and the accuracy of top-5 can be substantial[41, 47]. Situations such as those arising from the interpretation of chest X-rays show that, in some cases, the top-k predictions where $k>1$ hold significantly useful information when interpreting the predictions.
@@ 7 Conclusion and Future Work
This work has highlighted the importance of using holistic visual explanations for CNN predictions. It is observed that most of the existing saliency map generation methods, such as Grad-CAM solely focus on a single target class, which can neglect important information related to the reasoning behind CNN’s prediction. While investigating the related literature, it was identified that the substantial differences observed in top-1 and top-5 predictions can cause saliency maps to lose relevance when focusing only on the most probable model output.
To overcome the identified gaps in the existing methodologies, this work introduced FM-G-CAM, an unbiased methodology that generates saliency maps capable of providing a comprehensive interpretation of the CNN model decision-making process. FM-G-CAM can be configured to consider multiple top classes in the saliency map generation, providing a more holistic visual explanation for CNN predictions. FM-G-CAM is presented using detailed mathematical and algorithmic explanations, along with a practical implementation guide. Further, to highlight FM-G-CAM’s capabilities, its effectiveness was compared with that of Grad-CAM via real-world use cases. All code for the approach is released under an open-source licence.
For future work, the proposed approach uses gradient-weighted activations, though there are other approaches such as Grad-CAM++[11]. Toward this, future work could utilise other methods to provide further comparison to the proposed FM-G-CAM approach. One limitation of this study is within the approach to increase the visibility of lower activations, which promotes unwanted noise in some cases. Therefore, future work could explore methods to eliminate unwanted noise while retaining the visibility of lower-weighted neuroneal activations.
@@
To finally conclude, this work proposes a form of explainable artificial intelligence and contributes to the ongoing discussion surrounding more holistic approaches to explainability in image classification problems.

### v2

@@ 3.3 Choosing an optimal value for (K)
Unlike other saliency map generation methods, FM-G-CAM utilises multiple classes to generate the final saliency map. We recommend choosing the top predicted classes to be used. The top 3 or top 5 classes are recommended to be used with FM-G-CAM even though the algorithm allows any number of arbitrarily chosen classes to be used in the saliency map generation. However, this can be different and will depend mainly on below three factors: 1) The nature of the use case, 2) Average area of interest on the image for each predictive class, 3) Average number of distinct predictive labels per image.
@@ 3.3.1 Importance of L2 Normalisation
As shown in Figure 3, L2-normalisation across class gradient maps makes FM-G-CAM sensitive and unbiased, especially in the presence of classes with relatively lower weighted activations, while remaining accurate based on the content of the image. Also, as explained in the algorithm 1, normalisation is applied only after the filtering process. This allows filtering to be done based on raw gradient-weighted activations, independent of the normalisation process. L2-normalisation also helps to reduce unwanted noise, as shown in figure 2. This is because it is applied across the saliency maps by comparing the values against each other.
@@ 3.3.2 Choice of Colour Maps
Based on the findings presented in Section 2.3, we recommend using a single-colour colour map for each class when K in Section 3.3 is set to the recommended value of four. This will allow proper visualisation of the CNN decision-making process without overly complicating the saliency map. We comparatively test these in Section 6 with existing methods.
@@ 4 Implementation: XAI Inference Engine
For convenient use of FM-G-CAM, we made a Python library available to the community [^1] https://pypi.org/project/xai-inference-engine. The implementation is based on the algorithm 1. The code is open source and available on GitHub [^2] https://github.com/SuienS/xai-inference-engine. The library allows the user to use their trained PyTorch CNN-based models with our library. The library will output the prediction results along with the generated FM-G-CAM saliency map explaining the results. Figure 4 shows the overview of the XAI Inference Engine.
@@
The library also allows the end activation function used in the saliency map generation to be changed to ReLU, ELU, or GeLU. Figure 5 shows the comparative results of the use of the different activation functions available. It shows that GeLU and ELU tend to attract more noise than ReLU. However, at the same time, those can be useful when we want the saliency maps to be more sensitive. In contrast, Grad-CAM remains identical across all activation functions.
Refer to captionRefer to captionRefer to caption
@@
5 Quantitative Comparison: FM-G-CAM Vs Grad-CAM
With the intention of quantitatively evaluating the effectiveness of FM-G-CAM, we carried out a series of experiments that quantitatively compare FM-G-CAM with Grad-CAM. Only Grad-CAM was used in the experiments since we identified it as the theoretical basis even for later techniques such as Grad-CAM++ [11].
5.1 Methodology
Our tests uses four existing metrics for evaluating the reliability of a generated saliency map explaining a given vision-based model prediction. The metrics used in the tests are; Deletion Area Under Curve (DAUC) and Insertion Area Under Curve (IAUC) [41, 42] with Insertion Correlation (IC) and Deletion Correlation (DC) [43]. The DAUC metric is calculated by progressively masking the input image starting from the most important area on the image for a given class based on the provided saliency map. IAUC is also calculated similarly, but starts with a blank image and by revealing the most important areas sequentially until the entire image is revealed. DC and IC represent the correlation between the target class score and progressive masking or revealing steps similarly to DAUC and IAUC.
In section 2.3, we discussed the need for saliency maps to be based on multiple classes rather than a singular predicted top class. To objectively evaluate that FM-G-CAM provides a solution for the said issue by producing a fusion of multiple class activation maps, we repurposed above-mentioned metrics by calculating them repetitively on all top-5 predicted classes. With this method, we aimed to evaluate how these metrics varies with the ranks of top-5 predicted classes.
To carry out the planned tests, we used the ResNet50V2 [44] model with ImageNet weights [^3] https://pytorch.org/vision/main/models/generated/torchvision.models.resnet50.html and calculated the selected test metrics in the MS-COCO training set [45] with 80,000+ images. In summary, we used the knowledge of ImageNet to classify MS-COCO. MS-COCO was selected for this test since the majority of the images contain more than one ground truth class, which allows the planned test to be carried out meaningfully. For the calculation of these metrics we used an open source Python library [^4] https://github.com/TristanGomez44/metrics-saliency-maps introduced in [43]. We have also made our test scripts available open-source via GitHub [^5] https://github.com/SuienS/cam-evaluation.
5.2 Results
In this section, we present the results of the set of experiments as explained in the previous section. Table 3 presents the results of the tests carried out and for most test points, FM-G-CAM achieved higher scores than Grad-CAM. We also graphed this data as in Figure 6. The figure shows that both techniques perform similarly in metrics IAUC and DAUC, whereas FM-G-CAM consistently performs better in metrics IC and DC.
Table 3: Mean IAUC, IC, DAUC and DC of FM-G-CAM and Grad-CAM calculated for the top-5 predicted classes.
CAM Type Class Rank Mean IAUC Mean IC Mean DAUC Mean DC
FM-G-CAM 1 0.000345 -0.165391 0.000394 0.062597
2 0.001115 -0.156412 0.001328 0.080689
3 0.000330 -0.266839 0.000401 0.207409
4 0.000278 -0.260717 0.000285 0.185348
5 0.000283 -0.244950 0.000291 0.165731
Grad-CAM 1 0.000342 -0.184408 0.000394 0.056323
2 0.001119 -0.174904 0.001316 0.068046
3 0.000323 -0.278517 0.000397 0.196147
4 0.000277 -0.275418 0.000286 0.161246
5 0.000280 -0.269593 0.000292 0.151876
(a) IAUC
(b) DAUC
(c) IC
(d) DC
Figure 6: Results of the quantitative evaluation tests.
@@ 6 Use cases: A Comparative Practical Guide for FM-G-CAM
In this section, we compare our method with existing methods in generative saliency maps through a subset of use cases. It is worth noting that the method we propose is characteristically different from virtually all of the existing methods because of the nature of considering multiple classes for the saliency map. To be as fair as possible, we perform Grad-CAM on all the classes that FM-G-CAM uses to generate the saliency map, as can be seen in the following sections. Also, we purposefully compare our technique only with Grad-CAM, since any other improved or advanced version of Grad-CAM can be used within FM-G-CAM in the generation of the saliency map. We did not intend to improve the underlying theory of gradient-weighted activations for saliency map generation, but to utilise that idea in a way to provide a more holistic explanation of CNN predictions. The following sections describe some examples of FM-G-CAM use cases in comparison to Grad-CAM.
@@ 6.2 Medical AI
Another very interesting and important area where the proposed approach can be utilised is in medical AI. According to the literature, CNNs are a popular approach in the field of medical AI for a variety of diagnosis modalities [46, 47, 48, 49]. Therefore, an application of FM-G-CAM is explored in the diagnosis of chest X-Ray images. To this end, TorchXRayVision[50] weights are used for the DenseNet model[51]. Figure 8 shows the explained outputs when classifying the images selected from the validation set. The samples show that the FM-G-CAM presents a holistic and singular saliency map that can replace all other class-based Grad-CAM saliency maps. Furthermore, FM-G-CAM comparatively visualises the saliency maps of the selected classes, portraying their relative importance, whereas Grad-CAM considers each class in isolation, completely neglecting the fact that they are all part of the same prediction result.
@@ 7 Discussion
Using the CNN activations to explain their prediction was first introduced by [30], based on previous work investigating the capabilities of CNNs in object detection [52]. These works formed a basis for the idea of Grad-CAM [10]. However, these studies only consider a target class and, by doing so, disregard a considerable portion of information regarding the CNN classification process. In this study, we propose that the activations should be calculated for the top-$k$ classes, weighing them accordingly using their gradients, normalising them for a fair representation, and finally fusing them. Hence, while Grad-CAM and FM-G-CAM are based on the aforementioned studies, they are characteristically and mathematically different.
Figure 7 previously discussed shows that the single class Grad-CAM approach is not always useful in situations where a bigger-picture approach is needed. Instead, it only shows why a specified class has its related prediction score. In most cases, this class is selected as the top predicted class. However, there is a significant difference in the majority of CNN-based predictor models in the accuracy of top-1777https://paperswithcode.com/sota/image-classification-on-imagenet accuracy vs top-5888https://paperswithcode.com/sota/image-classification-on-imagenet?metric=Top%205%20Accuracy accuracy. The difference between the accuracy of top-1 and the accuracy of top-5 can be substantial[44, 51]. Situations such as those arising from the interpretation of chest X-rays show that, in some cases, the top-$k$ predictions where $k>1$ hold significantly useful information when interpreting the predictions.
@@ 8 Conclusion and Future Work
We highlighted the importance of using holistic visual explanations for CNN predictions in the previous section. It is observed that most of the existing saliency map generation methods, such as Grad-CAM solely focus on a single target class, which can neglect important information related to the reasoning behind CNN’s prediction. While investigating the related literature, we identified that the substantial differences observed in top-1 and top-5 predictions can cause saliency maps to lose relevance when focusing only on the most probable model output.
To overcome the identified gaps in existing methodologies, we introduced FM-G-CAM, an unbiased methodology that generates saliency maps capable of providing a comprehensive interpretation of the CNN model decision-making process. FM-G-CAM can be configured to consider multiple top classes in the saliency map generation, providing a more holistic visual explanation for CNN predictions. FM-G-CAM is presented using detailed mathematical and algorithmic explanations, along with a practical implementation guide. Furthermore, to highlight FM-G-CAM’s capabilities, we compared its effectiveness with that of Grad-CAM quantitatively via a customised test plan and qualitatively via real-world use cases. All the code for the approach is released under an open-source licence [^9] https://github.com/SuienS/cam-evaluation.
For future work, the proposed approach uses gradient-weighted activations, although there are other approaches such as Grad-CAM++[11]. To achieve this, future work could use other methods to provide a further comparison to the proposed FM-G-CAM approach. One limitation of our research is within the approach to increase the visibility of lower activations, which promotes unwanted noise in some cases. Therefore, future work could explore methods to eliminate unwanted noise while retaining the visibility of lower-weighted neuronal activations.
@@
To finally conclude, in this study, we propose a form of explainable artificial intelligence and contribute to the ongoing discussion surrounding more holistic approaches to explainability in image classification problems.